# Pushing JWST to the extremes: search and scrutiny of bright galaxy candidates at z15-30

### v1

@@ 5 Constraints on the UV Luminosity Function beyond z=15
We explore in this section the implications that our findings may have on the evolution of the Luminosity Function (LF) at . In order to take into account the caveats discussed above on the potential contamination from rare classes of interlopers, we consider two opposite scenarios regarding the reliability of our F200W-dropout sample. In Case 1, we assume that all our 9 candidates are at , as indicated by the selection function of Fig. 1. As we have described above, all our candidates are selected following a standard, self-consistent approach which has proved to be effective at lower redshifts. Despite the concerns described above, we stress that all our candidates are at least consistent with the expected properties of galaxies in this redshift range, and albeit contamination is expected both on the basis of our simulations and selection results (Sect. 3), there are no compelling reasons to reject them a priori. In the opposite Case 2 we assume that all candidates are interlopers, corresponding to a non-detection of sources in every field, and compute the corresponding upper limits on the LF.
@@ 5.1 The computation of the LF
The results are reported in Tables 2 (z=15-20) and 3 (z=20-30), and shown in Fig. 8 and 9, respectively.
The comparison to available measurements of the LF at clearly shows that if all our 9 F200W-dropout candidates are at (Case 1), the LF continues a trend of very slow evolution, similar to recent findings at (Harikane et al. 2025; Donnan et al. 2024; Robertson et al. 2024; Casey et al. 2024; Finkelstein et al. 2024). Our estimates would be in agreement with the measurements at z=16-18 by Harikane et al. (2024), Whitler et al. (2025), Kokorev et al. (2024) and Pérez-González et al. (2025) in the same luminosity range. Most importantly, a similar scenario remains valid if we consider partial, but not complete, contamination of our sample and/or successful confirmation of some candidates in our extended sample. In fact, every point in the LF is originated by a small number of observed galaxies falling in that luminosity bin (typically 1 or 2), such that even if a small fraction of candidates is indeed at , the corresponding density in the considered bin would be comparable to the available estimates at 14. In addition, three of our candidates are brighter than the highest-redshift secure galaxy JADES-GS-z14-0 (MUV=-20.81, Carniani et al. 2024a): if confirmed, they would imply a number density at M-21 higher than current measurements at 14.
@@
The estimate of the galaxy number density at , for which no reference F277W-dropout candidates have been selected, yields a similar conclusion. The derived upper limits are consistent with the z25 estimates by Pérez-González et al. (2025), and somewhat above the existing measurements at and at , due to the reduced effective volume sampled by our surveys. This result demonstrates that significantly wider/deeper areas are in principle necessary to sample this redshift range - a point that will be further expanded below.
@@ 5.2.2 Analytic and semi-analytic models
A number of self-consistent physical models have been explored to understand the high-abundance of bright galaxies observed at z9 by JWST. In the ?Attenuation-free model? (AFM, Ferrara et al. 2023; Ziparo et al. 2023; Fiore et al. 2023; Ferrara et al. 2025a) radiation-driven outflows expel or lift the previously formed dust, thus boosting the UV luminosity to an extent that matches the observed LFs at z10-14. In the ?Feedback-free starbursts? scenario (FFB, Dekel et al. 2023; Li et al. 2024) the excess of bright galaxies is explained as the result of high densities and low metallicities yielding a extremly high star-formation efficiency at cosmic dawn. The DELPHI semi-analytic model (SAM) based on cold gas fractions and star formation efficiencies sampled from the sphinx simulations (Mauerhofer et al. 2025) explored two different mechanisms boosting the abundance of galaxies at 9: a stellar initial mass function (IMF) that becomes increasingly top-heavy with decreasing metallicity and increasing redshift (eIMF model), and star formation efficiencies that increase with increasing redshift (eSFE model). Similarly, the CAT SAM invokes a gradual transition in the IMF, modulated by metallicity and redshift to match the UV LFs at very high-redshift (Trinca et al. 2024). Finally, the recently updated Santa Cruz semi-analytic model (Somerville et al. in prep., see also Yung et al. 2024a) was run on dark matter halo merger trees extracted from the GUREFT simulations (Yung et al. 2024b), and incorporates a star formation efficiency that increases with increasing gas surface density, motivated by results from molecular cloud-scale simulations with radiative transfer. As overall galaxy surface densities are naturally predicted to be higher at early times in the CDM picture, these models predict higher star formation efficiencies and therefore more UV luminous galaxies at early times. The free parameter represents the fraction of the ISM that is in dense, star forming clouds.
@@
Consistently with the simple empirical predictions described above, all these theoretical scenarios point towards a strong evolution at (Fig. 11). Both the AFM model, and the FFB one with star-formation efficiency , provide a good match to the z12 UV LF but predict an abundance of galaxies at z15 lower than our Case 1 LF. Similarly, the CAT SAM based on a standard IMF (Trinca et al. 2024), the DELPHI ?fiducial? model, and the Santa Cruz SAM with =0.1, fall below our estimates.
@@
Interestingly, all models are inconsistent at 2 with the Case 1 abundance at M-20, but the tension is alleviated when assuming a change in physical properties at z15. The DELPHI eIMF and eSFE models, and the CAT model assuming an evolving IMF, are partially consistent with our estimates at -18.5–19.5. An evolution of the dense gas fraction from 0.1 at z=13 to 0.5 could match our estimates at all luminosities according to the Santa Cruz SAM. The FFB model with is also consistent with our Case 1 results at -20, but lower SFEs may also be viable if galaxies at 15 follow a relation between stellar mass and UV luminosity different than the z10 one adopted by Li et al. (2024).
@@ 6 Designing a survey to break the z=15 barrier
As expected, we find that the bands immediately redward of the expected break are crucial for discriminating between low- and high-redshift solutions. Increasing the depth of the F090W and/or F115W bands is basically ineffective in reducing the contamination rate, while having both F150W and the F200W as deep as the LW bands reduces by a factor of 3 the fraction of templates contaminating the colour selections. The best advantage is obtained when the two bands are 0.5 mag deeper than the LW ones at fixed SNR. In such a case, the contamination fraction is reduced by a factor of 10. Most importantly, such a deeper imaging at 1.5-2 would make it possible to adopt the more inclusive color threshold (z15-20, see Fig. 1) with a contamination rate of 0.5% of the considered templates, with a significant gain in the accessible color space.
@@ 7 Summary and conclusions
We isolated nine candidates at , while no objects are found at . A closer inspection of the selected candidates shows that despite exhibiting a 1.5 mag break, the selected objects consistently display multimodal redshift probability distributions across different SED-fitting codes and methodologies. The alternative solutions cover regions in the versus and versus planes different from the general populations at the same redshifts. Most importantly, they correspond to populations of low-mass () quiescent or dusty galaxies that have not been thoroughly investigated so far. This result is corroborated by the spectral properties of object UDS_56824, a confirmed interloper of our F200W-dropout selection, which is found to be a dusty () starburst galaxy at with mass . These results imply that while our candidates are, in principle, credible objects at , none of them would pass stringent selection criteria based on between different redshift solutions. In addition, considering that the low-redshift templates populate a basically unexplored parameter space whose galaxy density is unconstrained, using their as a weight to estimate the high-redshift UV LF appears to be an approach rife with uncertainties.
We adopted a more pragmatic approach of estimating the UV LF at z15-20 by assuming different contamination levels. The UV LF based on the, admittedly extreme, assumption of negligible contamination indicates a very mild evolution compared to estimates at immediately lower redshifts, at odds with all theoretical predictions. In particular, the confirmation of bright (M-21) candidates would require deep revisions of our theoretical framework, and might be even in contrast with any plausible model under a standard –CDM cosmological scenario. However, the tension with theoretical models is so significant that even if only a small fraction of the candidates is confirmed to be at , a further evolution of physical properties, such as IMF or star-formation efficiency would be required to explain the observed number densities. If instead all the analysed candidates are interlopers, we are forced to conclude that future surveys will need to cover much wider areas to secure the selection of bright galaxies significantly beyond .
@@ Acknowledgements.
We thank J. Dunlop and R. Ellis for the useful discussions. We acknowledge financial support from NASA through grant JWST-ERS-1324. Support was also provided by the PRIN 2022 MUR project 2022CB3PJ3 – First Light And Galaxy aSsembly (FLAGS) funded by the European Union – Next Generation EU, and by INAF Mini-grant “Reionization and Fundamental Cosmology with High-Redshift Galaxies.” P.G.P.-G. and L.C. acknowledge support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa. L.N. acknowledges support from grant “Progetti per Avvio alla Ricerca - Tipo 1, Unveiling Cosmic Dawn: Galaxy Evolution with CAPERS” (AR1241906F947685). RA acknowledges financial support from grants PID2023-147386NB-I00 ”XTREM” and PID2022-136598NB-C32 “Estallidos8”, funded by MCIN/AEI/10.13039/501100011033 and ERDF/EU, and the State Agency for Research of the Spanish MCIU through ‘Center of Excellence Severo Ochoa’ award to the IAA-CSIC (SEV-2017-0709) and CEX2021-001131-S. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. The Flatiron Institute is supported by the Simons Foundation. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1324, 1345, 1180, 1210, 1837, 2561, 2756, 3990, and 6368.References
@@
- Bergamini et al. (2023) Bergamini, P., Acebron, A., Grillo, C., et al. 2023, arXiv e-prints, arXiv:2303.10210
@@
- Bezanson et al. (2022) Bezanson, R., Labbe, I., Whitaker, K. E., et al. 2022, arXiv e-prints, arXiv:2212.04026
@@
- Carniani et al. (2024a) Carniani, S., D’Eugenio, F., Ji, X., et al. 2024a, arXiv e-prints, arXiv:2409.20533
- Carniani et al. (2024b) Carniani, S., Hainline, K., D’Eugenio, F., et al. 2024b, Nature, 633, 318
@@
- Conselice et al. (2024) Conselice, C. J., Adams, N., Harvey, T., et al. 2024, arXiv e-prints, arXiv:2407.14973
@@
- Finkelstein et al. (2025) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2025, arXiv e-prints, arXiv:2501.04085
@@
- Finkelstein et al. (2023) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2023, arXiv e-prints, arXiv:2311.04279
- Finkelstein et al. (2024) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024, ApJ, 969, L2
@@
- Kokorev et al. (2024) Kokorev, V., Atek, H., Chisholm, J., et al. 2024, arXiv e-prints, arXiv:2411.13640
@@
- Larson et al. (2022) Larson, R. L., Hutchison, T. A., Bagley, M., et al. 2022, arXiv e-prints, arXiv:2211.10035
@@
- Mauerhofer et al. (2025) Mauerhofer, V., Dayal, P., Haehnelt, M. G., et al. 2025, arXiv e-prints, arXiv:2502.02647
@@
- Merlin et al. (2022) Merlin, E., Bonchi, A., Paris, D., et al. 2022, arXiv e-prints, arXiv:2207.11701
@@
- Morishita et al. (2024) Morishita, T., Mason, C. A., Kreilgaard, K. C., et al. 2024, arXiv e-prints, arXiv:2412.04211
@@
- Pérez-González et al. (2023b) Pérez-González, P. G., Costantin, L., Langeroodi, D., et al. 2023b, arXiv e-prints, arXiv:2302.02429
- Pérez-González et al. (2024b) Pérez-González, P. G., D‘Eugenio, F., Rodríguez del Pino, B., et al. 2024b, arXiv e-prints, arXiv:2405.03744
- Pérez-González et al. (2025) Pérez-González, P. G., Östlin, G., Costantin, L., et al. 2025, arXiv e-prints, arXiv:2503.15594
- Pérez-González et al. (2024c) Pérez-González, P. G., Rinaldi, P., Caputi, K. I., et al. 2024c, ApJ, 969, L10
@@
- Roberts-Borsani et al. (2024a) Roberts-Borsani, G., Bagley, M., Rojas-Ruiz, S., et al. 2024a, arXiv e-prints, arXiv:2407.17551
- Roberts-Borsani et al. (2024b) Roberts-Borsani, G., Treu, T., Shapley, A., et al. 2024b, arXiv e-prints, arXiv:2403.07103
@@
- Tacchella et al. (2023) Tacchella, S., Eisenstein, D. J., Hainline, K., et al. 2023, ApJ, 952, 74
@@
- Weibel et al. (2024) Weibel, A., de Graaff, A., Setton, D. J., et al. 2024, arXiv e-prints, arXiv:2409.03829
- Westcott et al. (2024) Westcott, L., Conselice, C. J., Harvey, T., et al. 2024, arXiv e-prints, arXiv:2412.14970

### v2

@@ 5 Constraints on the UV Luminosity Function beyond z=15
We explore in this section the implications that our findings may have on the evolution of the Luminosity Function (LF) at . In order to take into account the caveats discussed above on the potential contamination from rare classes of interlopers, we consider two opposite scenarios regarding the reliability of our F200W-dropout sample. In Case 1, we assume that all our five candidates are at , as indicated by the selection function of Fig. 1. As we have described above, all our candidates are selected following a standard, self-consistent approach which has proved to be effective at lower redshifts. Despite the concerns described above, we stress that all our candidates are at least consistent with the expected properties of galaxies in this redshift range, and albeit contamination is expected both on the basis of our simulations and selection results (Sect. 3), there are no compelling reasons to reject them a priori. In the opposite Case 2 we assume that all candidates are interlopers, corresponding to a non-detection of sources in every field, and compute the corresponding upper limits on the LF.
@@ 5.1 The computation of the LF
The results are reported in Tables 2 (z=15-20) and 3 (z=20-30), and shown in Fig. 7 and 8, respectively.
The comparison to available measurements of the LF at clearly shows that if all our 5 F200W-dropout candidates are at (Case 1), the LF continues a trend of very slow evolution, similar to recent findings at (Harikane et al. 2025; Donnan et al. 2024; Robertson et al. 2024; Casey et al. 2024; Finkelstein et al. 2024b). Our estimates would be in agreement with the measurements at z=16-18 by Harikane et al. (2024), Whitler et al. (2025), Kokorev et al. (2025) Pérez-González et al. (2025b), and Weibel et al. (2025b) in the same luminosity range. Most importantly, a similar scenario remains valid if we consider partial, but not complete, contamination of our sample and/or successful confirmation of some candidates in our extended sample. In fact, every point in the LF is originated by a small number of observed galaxies falling in that luminosity bin (typically 1 or 2), such that even if a small fraction of candidates is indeed at , the corresponding density in the considered bin would be comparable to the available estimates at 14. In addition, two of our candidates are brighter than the highest-redshift secure galaxy JADES-GS-z14-0 (MUV=-20.81, Carniani et al. 2025): if confirmed, they would imply a number density at M-21 higher than current measurements at 14.
@@
The estimate of the galaxy number density at , for which no reference F277W-dropout candidates have been selected, yields a similar conclusion. The derived upper limits are consistent with the z25 estimates by Pérez-González et al. (2025b), and somewhat above the existing measurements at and at , due to the reduced effective volume sampled by our surveys. This result demonstrates that significantly wider/deeper areas are in principle necessary to sample this redshift range - a point that will be further expanded below.
@@ 5.2.2 Analytic and semi-analytic models
A number of self-consistent physical models have been explored to understand the high-abundance of bright galaxies observed at z9 by JWST. In the ?Attenuation-free model? (AFM, Ferrara et al. 2023; Ziparo et al. 2023; Fiore et al. 2023; Ferrara et al. 2025a) radiation-driven outflows expel or lift the previously formed dust, thus boosting the UV luminosity to an extent that matches the observed LFs at z10-14. In the ?Feedback-free starbursts? scenario (FFB, Dekel et al. 2023; Li et al. 2024) the excess of bright galaxies is explained as the result of high densities and low metallicities yielding a extremly high star-formation efficiency at cosmic dawn444The predictions used in the present work incorporate the improved estimate of the UV luminosity based on the median SFH, as described in Dekel et al. (2025).. The DELPHI semi-analytic model (SAM) based on cold gas fractions and star formation efficiencies sampled from the sphinx simulations (Mauerhofer et al. 2025) explored two different mechanisms boosting the abundance of galaxies at 9: a stellar initial mass function (IMF) that becomes increasingly top-heavy with decreasing metallicity and increasing redshift (eIMF model), and star formation efficiencies that increase with increasing redshift (eSFE model). Similarly, the CAT SAM invokes a gradual transition in the IMF, modulated by metallicity and redshift to match the UV LFs at very high-redshift (Trinca et al. 2024). Finally, the recently updated Santa Cruz semi-analytic model (Yung et al. 2024a; Somerville et al. 2025) was run on dark matter halo merger trees extracted from the GUREFT simulations (Yung et al. 2024b), and incorporates a star formation efficiency that increases with increasing gas surface density, motivated by results from molecular cloud-scale simulations with radiative transfer. As overall galaxy surface densities are naturally predicted to be higher at early times in the CDM picture, these models predict higher star formation efficiencies and therefore more UV luminous galaxies at early times. The free parameter represents the fraction of the ISM that is in dense, star forming clouds.
@@
Consistently with the simple empirical predictions described above, all these theoretical scenarios point towards a strong evolution at (Fig. 10). Both the AFM model, and the FFB one with star-formation efficiency , provide a good match to the z12 UV LF but while the AFM predicts an abundance of galaxies at z15 lower than our Case 1 LF, the FFB model with is consistent with our estimates at M-22. The CAT SAM based on a standard IMF (Trinca et al. 2024), the DELPHI ?fiducial? model, and the Santa Cruz SAM with =0.1, fall below our estimates.
@@
Interestingly, 2 tensions are found at the brightest magnitudes, i.e. at M-22 in the case of the FFB model, and at M-20 for the others. The tensions are apparently alleviated when assuming a change in physical properties. The DELPHI eIMF and eSFE models, and the CAT model assuming an evolving IMF, are partially consistent with our estimates at -18.5–19.5. An evolution of the dense gas fraction from 0.1 at z=13 to 0.5 could match our estimates at all luminosities according to the Santa Cruz SAM, while the FFB model with would match the observed Case 1 abundance in the brightest bin.
@@ 6 Designing a survey to break the z=15 barrier
As expected, we find that the bands immediately redward of the expected break are crucial for discriminating between low- and high-redshift solutions. Increasing the depth of the F090W and/or F115W bands is basically ineffective in reducing the contamination rate, while having both F150W and the F200W as deep as the LW bands reduces by a factor of 3 the fraction of templates contaminating the colour selections. The best advantage is obtained when the two bands are 0.5 mag deeper than the LW ones at fixed SNR. In such a case, the contamination fraction is reduced by a factor of 10. Most importantly, such a deeper imaging at 1.5-2 would make it possible to adopt the more inclusive color threshold (z15-20, see Fig. 1) with a contamination rate of 0.5% of the considered templates, with a significant gain in the accessible color space. Our analysis is in good agreement with the results by Adams et al. (2025), showing that doubling the exposure time in the bands sampling below the break is the most efficient way to reduce contamination in a standard high-redshift survey.
@@ 7 Summary and conclusions
We isolated five candidates at , while no objects are found at . A closer inspection of the selected candidates shows that despite exhibiting a 1.5 mag break, the selected objects consistently display multimodal redshift probability distributions across different SED-fitting codes and methodologies. The alternative solutions cover regions in the versus and versus planes different from the general populations at the same redshifts. Most importantly, they correspond to populations of low-mass () quiescent or dusty galaxies that have not been thoroughly investigated so far. This result is corroborated by the spectral properties of five confirmed interlopers of our F200W-dropout selection. Four objects are found to be dusty star-forming galaxies, including three low-mass (7.6-8.8) objects at z2-5, and a highly attenuated () starburst galaxy at with mass . The fifth confirmed interloper, CEERS_15937 at z=4.91, is found to be the first spectroscopically confirmed passive galaxy with log 9 at high-redshift. These results imply that while our candidates are, in principle, credible objects at , none of them would pass stringent selection criteria based on between different redshift solutions. In addition, considering that the low-redshift templates populate a basically unexplored parameter space whose galaxy density is unconstrained, using their as a weight to estimate the high-redshift UV LF appears to be an approach rife with uncertainties.
We adopted a more pragmatic approach of estimating the UV LF at z15-20 by assuming different contamination levels. The UV LF based on the, admittedly extreme, assumption of negligible contamination indicates a very mild evolution compared to estimates at immediately lower redshifts, at odds with all theoretical predictions. In particular, the confirmation of bright (M-21) candidates would require deep revisions of our theoretical framework, and might be even in contrast with any plausible model under a standard –CDM cosmological scenario. The tension with theoretical models implies that even if only a small fraction of the candidates is confirmed to be at , a further evolution of physical properties, such as IMF or star-formation efficiency would be required to explain the observed number densities. If instead all the analysed candidates are interlopers, we are forced to conclude that future surveys will need to cover much wider areas to secure the selection of bright galaxies significantly beyond .
@@ Acknowledgements.
We thank the referee for the constructive comments that helped us improve the manuscript. We thank J. Dunlop and R. Ellis for the useful discussions. We acknowledge financial support from NASA through grant JWST-ERS-1324. Support was also provided by the PRIN 2022 MUR project 2022CB3PJ3 – First Light And Galaxy aSsembly (FLAGS) funded by the European Union – Next Generation EU, by INAF Mini-grant “Reionization and Fundamental Cosmology with High-Redshift Galaxies”, and by INAF GO Grant ”Revealing the nature of bright galaxies at cosmic dawn with deep JWST spectroscopy.” P.G.P.-G. and L.C. acknowledge support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa. L.N. acknowledges support from grant “Progetti per Avvio alla Ricerca - Tipo 1, Unveiling Cosmic Dawn: Galaxy Evolution with CAPERS” (AR1241906F947685). RA acknowledges financial support from grants PID2023-147386NB-I00 ”XTREM” and PID2022-136598NB-C32 “Estallidos8”, funded by MCIU/AEI/10.13039/501100011033 and ERDF/EU, and the State Agency for Research of the Spanish MCIU through ‘Center of Excellence Severo Ochoa’ award to the IAA-CSIC (SEV-2017-0709) and CEX2021-001131-S. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. The Flatiron Institute is supported by the Simons Foundation. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1324, 1345, 1180, 1210, 1837, 2561, 2756, 3794, 3990, 5893, and 6368. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2015.1.01074.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. Some of the data products presented herein were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.References
- Adams et al. (2025) Adams, N. J., Austin, D., Harvey, T., et al. 2025, MNRAS, 542, 1705
@@
- Baker et al. (2025) Baker, W. M., Ito, K., Valentino, F., et al. 2025, arXiv e-prints, arXiv:2509.09761
- Bergamini et al. (2023) Bergamini, P., Acebron, A., Grillo, C., et al. 2023, ApJ, 952, 84
@@
- Bezanson et al. (2024) Bezanson, R., Labbe, I., Whitaker, K. E., et al. 2024, ApJ, 974, 92
@@
- Carniani et al. (2025) Carniani, S., D’Eugenio, F., Ji, X., et al. 2025, A&A, 696, A87
- Carniani et al. (2024) Carniani, S., Hainline, K., D’Eugenio, F., et al. 2024, Nature, 633, 318
@@
- Conselice et al. (2025) Conselice, C. J., Adams, N., Harvey, T., et al. 2025, ApJ, 983, 30
@@
- Dekel et al. (2025) Dekel, A., Mandelker, N., Li, Z., et al. 2025, arXiv e-prints, arXiv:2506.11664
@@
- Finkelstein et al. (2025) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2025, ApJ, 983, L4
@@
- Finkelstein et al. (2024a) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024a, ApJ, 969, L2
- Finkelstein et al. (2024b) Finkelstein, S. L., Leung, G. C. K., Bagley, M. B., et al. 2024b, ApJ, 969, L2
@@
- Kokorev et al. (2025) Kokorev, V., Atek, H., Chisholm, J., et al. 2025, ApJ, 983, L22
@@
- Larson et al. (2023) Larson, R. L., Hutchison, T. A., Bagley, M., et al. 2023, ApJ, 958, 141
@@
- Mauerhofer et al. (2025) Mauerhofer, V., Dayal, P., Haehnelt, M. G., et al. 2025, A&A, 696, A157
@@
- Merlin et al. (2022) Merlin, E., Bonchi, A., Paris, D., et al. 2022, ApJ, 938, L14
- Merlin et al. (2025) Merlin, E., Fortuni, F., Calabró, A., et al. 2025, arXiv e-prints, arXiv:2509.09764
@@
- Morishita et al. (2025) Morishita, T., Mason, C. A., Kreilgaard, K. C., et al. 2025, ApJ, 983, 152
@@
- Naidu et al. (2025) Naidu, R. P., Oesch, P. A., Brammer, G., et al. 2025, arXiv e-prints, arXiv:2505.11263
@@
- Pérez-González et al. (2023b) Pérez-González, P. G., Costantin, L., Langeroodi, D., et al. 2023b, ApJ, 951, L1
- Pérez-González et al. (2025a) Pérez-González, P. G., D’Eugenio, F., Rodríguez del Pino, B., et al. 2025a, Nature Astronomy, 9, 1240
- Pérez-González et al. (2025b) Pérez-González, P. G., Östlin, G., Costantin, L., et al. 2025b, ApJ, 991, 179
- Pérez-González et al. (2024b) Pérez-González, P. G., Rinaldi, P., Caputi, K. I., et al. 2024b, ApJ, 969, L10
@@
- Roberts-Borsani et al. (2025a) Roberts-Borsani, G., Bagley, M., Rojas-Ruiz, S., et al. 2025a, ApJ, 983, 18
- Roberts-Borsani et al. (2025b) Roberts-Borsani, G., Oesch, P., Ellis, R., et al. 2025b, arXiv e-prints, arXiv:2508.21708
- Roberts-Borsani et al. (2024) Roberts-Borsani, G., Treu, T., Shapley, A., et al. 2024, ApJ, 976, 193
@@
- Somerville et al. (2025) Somerville, R. S., Yung, L. Y. A., Lancaster, L., et al. 2025, arXiv e-prints, arXiv:2505.05442
@@
- Strait et al. (2023) Strait, V., Brammer, G., Muzzin, A., et al. 2023, ApJ, 949, L23
@@
- Weibel et al. (2025a) Weibel, A., de Graaff, A., Setton, D. J., et al. 2025a, ApJ, 983, 11
- Weibel et al. (2025b) Weibel, A., Oesch, P. A., Williams, C. C., et al. 2025b, arXiv e-prints, arXiv:2507.06292
- Westcott et al. (2025) Westcott, L., Conselice, C. J., Harvey, T., et al. 2025, ApJ, 983, 121